# Self-Ensembling Gaussian Splatting for Few-shot Novel View Synthesis

### v1

@@ 3.2 Temporal Samples in 3DGS
Figure 4: Buffer update during training. For each sampled pseudo view, we dynamically update the buffer which stores the images rendered at different training steps. For instance, at training step $t_{T}$, the oldest image $\mathbf{I}_{T-S}$ in the buffer is popped, and the new image $\mathbf{I}_{T}$ is pushed into the buffer. An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed for the perturbation.
@@
In contrast to generating multiple samples by training separate 3DGS models, we dynamically create temporal samples from a single $\mathbf{\Delta}$-model during training. Specifically, we train the $\mathbf{\Delta}$-model on the available posed images, following the same optimization and density control strategies introduced in 3DGS [20]. At each training iteration, the current $\mathbf{\Delta}$-model represents a specific sample in the Gaussian parameter space based on the information contained in the training data. A temporal 3DGS model is then sampled from the neighboring region of the $\mathbf{\Delta}$-model in the parameter space by perturbing the $\mathbf{\Delta}$-model as:
@@
where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at the training step $t$, $\delta_{t}$ indicates the perturbations, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the temporal model. To achieve this, a plausible approach is to sample random noise from a normal distribution, $\delta_{t}\in\mathcal{N}(\mu_{t},\sigma_{t})$. However, since these perturbations are independent of the training data, this method tends to result in invalid samples.
Consequently, we present an uncertainty-aware perturbing strategy according to the statistics on the renderings. The approach starts by sampling pseudo views along the camera trajectories of the training views. Given two cameras sampled from the training views, the camera extrinsic of a pseudo view is computed as:
@@
where $(\mathbf{R}_{1},\mathbf{R}_{2})$ represent the rotations of the sampled cameras, $(\mathbf{c}_{1},\mathbf{c}_{2})$ indicate the sampled camera centers, SLERP denotes the spherical linear interpolation [4], $\beta$ is a scalar that controls the interpolation, and $(\hat{\mathbf{R}},\hat{\mathbf{T}})$ stand for the camera parameters of the pseudo view. As illustrated in Fig. 3, for each pseudo view, we render an RGB image using the current $\mathbf{\Delta}$-model and store the renderings at different training steps in a buffer. The uncertainties of these renderings are then computed as:
$\mathbf{U}=\frac{1}{S}\sqrt{\sum_{i=1}^{S}(\mathbf{I}_{i}-\bar{\mathbf{I}})^{2}},$ (8)
where $\mathbf{U}\in\mathbb{R}^{H\times W}$ is the uncertainty map with the same resolution as the rendered image, $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size. This uncertainty estimation is carried out over all sampled pseudo views in parallel. In practice, to enhance the robustness, we perform local smoothing over each uncertainty map, which is denoted as:
@@
where $k=5$ is the kernel size. Subsequently, we apply the uncertainty-aware perturbations to each of the Gaussians in the $\mathbf{\Delta}$-model, formulated as:
@@
where $\hat{\mathcal{U}}^{t}$ is a set of the computed uncertainty maps at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as:
@@
in which $\tau$ is a predefined threshold and $u_{ij}$ is the uncertainty score of a pixel sampled from a set of pixels $\mathcal{P}$ overlapping with the 2D splat of $G_{\Delta}^{t}$. Please refer to the supplementary material for details on defining $\tau$ and identifying $\mathcal{P}$. Therefore, the reliability of the Gaussians is connected with the uncertainties of renderings, and only the unreliable Gaussians characterized by high uncertainty scores are perturbed. In practice, we add the perturbations to the 3D points, rotations, scales, and opacities of the $\mathbf{\Delta}$-model. Particularly, since the 3D rotation is not contiguous when formulated as a 3D matrix, we perturb the 6D continuous representation [51] which is defined as:
$\hat{\mathbf{R}}^{t}_{\Delta}=f^{-1}(f(\mathbf{R}_{\Delta}^{t})+\delta^{R}_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})),\ \ \delta^{R}_{t}\in\mathbb{R}^{6}.$ (12)
Fig. 4 illustrates the dynamic updates of the buffer throughout the training process. As the buffer is updated, the uncertainty map varies accordingly, resulting in diverse temporal samples from the $\mathbf{\Delta}$-model.
Moreover, it is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to directly compute the uncertainties of the Gaussian parameters across different training steps. The proposed perturbing approach can handle this issue since the statistics on 2D renderings provide a consistent and efficient assessment, regardless of the varying numbers of Gaussians.
@@ 3.3 Self-Ensembling in 3DGS
Given the temporal samples around the $\mathbf{\Delta}$-model, the next step in our pipeline is to construct an ensemble from these samples. Concretely, we train a separate 3DGS model named the $\mathbf{\Sigma}$-model on the posed images without perturbations. To prevent the $\mathbf{\Sigma}$-model from being stuck at local optima, we guide the training of the $\mathbf{\Sigma}$-model by leveraging a regularization technique [48]. The regularization is formulated as:
@@
where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term, and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and $\mathbf{\Delta}$-model, respectively. We also utilize a co-pruning strategy [48] to enhance the regularization. Therefore, the $\mathbf{\Sigma}$-model aggregates information from different samples in the parameter space, serving as an ensemble of 3DGS models. The final loss function during training is defined as:
@@
where $\gamma=1$ by default and $\mathcal{L}_{\text{RGB}}$ represents a photometric loss over the training views. At testing, we keep the $\mathbf{\Sigma}$-model for novel view synthesis. Notably, in contrast to CoR-GS, we derive these samples from a single $\mathbf{\Delta}$-model without additional training costs. As a result, our SE-GS can be efficiently scaled up to a large number of 3DGS models, enabling more effective regularization. Moreover, our regularization is carried out on pseudo views in a self-supervised manner, which is thus independent of additional ground truth such as monocular depth [23, 53].
@@ 4.1 Setup
Datasets. We conduct experiments on four datasets, i.e., LLFF [25], DTU [19], Mip-NeRF360 [3], and MVImgNet [46]. On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in CoR-GS [48], using the same training/testing splits and downsampling rates. As suggested in CoR-GS, we mask the background when assessing the quality of novel view synthesis on the DTU dataset. Since LLFF, DTU, and Mip-NeRF360 offer a limited number of scenarios, we extend our experiment to a large-scale dataset, i.e., MVImgNet. We randomly sample 50 scenes from this dataset and resize the longest side of each image to 512. The Gaussian parameters for all evaluated methods are initialized based on the point clouds obtained from COLMAP [34] to ensure a fair comparison. It is worth noting that COLMAP fails on certain scenes within the DTU dataset in the setting of sparse views, where we instead randomly initialize the point clouds. Please refer to the supplementary material for more details on the setup.
Implementation details. We train our SE-GS for 10,000 iterations on the LLFF, DTU, and MVImgNet datasets, and for 30,000 iterations on Mip-NeRF360. In the uncertainty-aware perturbing mechanism, we employ 24 image buffers with a buffer size of 3 and perturb the $\mathbf{\Delta}$-model with an interval of 500 iterations. The noise $\delta_{t}$ in Eq. 10 is sampled from a standard normal distribution, i.e., $\delta_{t}\in\mathcal{N}(0,1)$. More details are provided in the supplementary material.
@@ 4.2 Quantitative results
We list the quantitative results of the evaluated approaches on the LLFF, DTU, Mip-NeRF360, and MVImgNet datasets in Table 1, Table 2, Table 3, and Table 4, respectively. The best, second-best, and third-best results in each column are highlighted in red, orange, and yellow, respectively. For MVSplat [7], we employ the pretrained model released by the authors and evaluate it on DTU. Notably, the per-scene optimized 3DGS methods significantly outperform MVSplat, highlighting the advantages of scene-level radiance field modeling for high-quality novel view synthesis. The performance of both NeRF and 3DGS methods consistently declines as the number of training views decreases, underscoring the challenge of few-shot novel view synthesis. In this context, our method surpasses all competitors across the four datasets. Note that in the scenario of sparse views, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
@@
Specifically, on the LLFF and DTU datasets, we train the methods using 3, 6, and 9 views. Our method achieves an improvement of $0.34$ in terms of PSNR with 3 views on LLFF. The improvement is more significant on DTU, reaching a gain of $0.59$. Since the images in the Mip-NeRF360 dataset are captured from diverse viewpoints, we sample more training views, i.e., 12 and 24, following the setting used in CoR-GS. On this dataset, our SE-GS achieves the highest PSNR and SSIM, as well as the lowest LPIPS and AVGE, demonstrating superior NVS quality under the challenges of sparse views with large-scale viewpoint variations. Additionally, to further assess the robustness across diverse scenarios, we conduct an experiment on 50 scenes sampled from the MVImgNet dataset. As reported in Table 4, our method obtains consistently better results than previous state-of-the-art approaches when trained with 5, 7, and 10 views. Our SE-GS also demonstrates better compatibility with relatively dense views. For instance, the PSNR improvement increases from $0.26$ to $0.59$ as the number of training views varies from 5 to 10.
@@
Methods like DNGaussian [23] and FSGS [53], which leverage auxiliary data terms sometimes outperform 3DGS. However, in some cases, such as 9 views on the LLFF dataset, the PSNR of FSGS becomes worse than 3DGS. Since the data acquired through off-the-shelf approaches might be unreliable, the quality of the NVS results is consequently affected. Compared with FSGS and DNGaussian, CoR-GS [48] exhibits better stability, outperforming 3DGS in more scenarios. This finding highlights the promising potential of regularization in the sparse-view setting. However, the improvement is still not consistent, as CoR-GS shows worse LPIPS than 3DGS when trained with 10 views on the MVImgNet dataset. In contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability. Our method is capable of avoiding specific local optima since the $\mathbf{\Sigma}$-model aggregates information from diverse temporal samples in the presented self-ensembling mechanism.
4.3 Qualitative results.
@@
Fig. 5 presents qualitative comparisons among 3DGS, CoR-GS, and our SE-GS, illustrating renderings from novel views. The models are trained on the DTU and MVImgNet datasets with 3 and 5 posed images, respectively. Our SE-GS produces fewer visual artifacts compared to the other methods, achieving better robustness against the challenge of few-shot training views. Moreover, our method captures finer details, particularly in areas with complex and repeated textures such as the pineapple and watch strap. This demonstrates the effectiveness of our self-ensembling strategy in improving both the visual quality and stability of novel view neural rendering under limited-view conditions.
@@
Figure 6: Comparison with CoR-GS. CoR-GS [48] is trained with different numbers of Gaussian models. The number varies from 2 to 8. PSNR on LLFF with 3 training views is utilized as the metric to evaluate the NVS quality. Iterations/s refers to the number of training iterations completed per second, indicating the training speed.
@@ 4.4 Analysis
To shed more light on the effectiveness and efficiency of our SE-GS, we develop a comprehensive analysis of the introduced self-ensembling mechanism. The analysis starts by comparing our approach with CoR-GS trained with varying numbers of Gaussian models. The detailed results on LLFF with 3 training views are shown in Fig. 6. As illustrated in Fig. 6(a), the PNSR of CoR-GS increases as more Gaussian models are trained. Since adding more Gaussian models enhances the regularization effect in CoR-GS, this observation highlights the potential of improving 3DGS in the sparse-view setting by strengthening the regularization process. The PSNR reaches the peak with 6 Gaussian models, indicating an upper bound when using a limited number of models. In contrast, our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in much more diverse samples of Gaussian models compared to CoR-GS. This leads to more effective regularization, as reflected by the better PSNR score.
Moreover, Fig. 6(b) shows the training speed measured by the number of training iterations completed per second. As witnessed in this figure, the speed of CoR-GS significantly drops as the number of Gaussian models increases. This limitation poses a challenge in scaling CoR-GS up to a large number of Gaussian models. Note that in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model. The temporal samples are generated through perturbations. The result in Fig. 6(b) shows that we achieve a comparable training speed to CoR-GS with 2 models. Only a negligible additional cost is incurred due to the perturbation process. It is worth noting that only the $\mathbf{\Sigma}$-model is retained for neural rendering during inference, so the testing speed of SE-GS is the same as that of the original 3DGS. Consequently, our method is capable of efficiently enhancing the performance of 3DGS with sparse training views.
@@
Table 5: Effectiveness of the uncertainty-aware perturbations. Random means we add random perturbations to all Gaussians in the $\mathbf{\Delta}$-model. Gradient-Aware indicates an alternative in which we perturb the Gaussians based on gradients. All methods are trained on LLFF with 3 views for 2000 iterations and the PSNR is utilized as the metric.
Furthermore, we assess the effectiveness of our uncertainty-aware perturbing strategy, comparing it with two alternatives. Specifically, as we introduced in Sec. 3, the straightforward method perturbing the $\mathbf{\Delta}$-model is to add random noise to all Gaussians in the model. We denote this approach as Random. Building upon this baseline, an alternative approach, referred to as Gradient-Aware, identifies unreliable Gaussians by analyzing their gradients. Specifically, we replace the indicator function in Eq. 11 with one based on gradient magnitude. Gaussians with gradients exceeding a threshold are perturbed. In practice, we adopt the same threshold as used in density control. We train these two alternatives and our approaches on LLFF with 3 views for 2000 iterations, which challenges the methods in terms of convergence speed. As listed in Table 5, Random performs the worst since it leads to many invalid Gaussian samples, negatively impacting convergence speed. Note that our uncertainty-aware strategy aggregates information from multiple training steps, while the gradient-aware method can only utilize the gradients at the current training step. The improvement in PSNR shown in Table 5 demonstrates the superior effectiveness of our approach.

### v2

@@ 3.2 Uncertainty-Aware Perturbation
Figure 4: Buffer update during training. For each sampled pseudo view, we dynamically update the buffer that stores the images rendered at different training steps. For instance, at training step $t_{T}$, the oldest image $\mathbf{I}_{T-S}$ in the buffer is popped, and the new image $\mathbf{I}_{T}$ is pushed into the buffer. An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed to compute a perturbation that results in a new 3DGS model.
@@
In contrast to generating multiple samples by training separate 3DGS models, we dynamically create temporal samples from a single $\mathbf{\Delta}$-model during training. Note that, in this paper, temporal refers to the dynamic nature of samples that vary across training iterations. Specifically, we train the $\mathbf{\Delta}$-model on the available posed images, following the same optimization and density control strategies introduced in 3DGS [21]. At each training iteration, the current $\mathbf{\Delta}$-model represents a specific sample in the Gaussian parameter space based on the information contained in the training data. A temporal 3DGS model is then sampled from the neighboring region of the current $\mathbf{\Delta}$-model in the parameter space by perturbing the $\mathbf{\Delta}$-model as
@@
where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at training step $t$, $\delta_{t}$ indicates the perturbation, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the temporal model. To achieve the perturbation, a naive approach is to sample random noise from a normal distribution, i.e., $\delta_{t}\in\mathcal{N}(\mu_{t},\sigma_{t}^{2})$. However, such perturbation is independent of the training data, and this method thus tends to yield invalid samples.
Therefore, we present an uncertainty-aware perturbation strategy, leveraging the statistics of renderings. Our approach starts by creating $M$ pseudo views through interpolation between the training views. Specifically, given two cameras sampled from the training views, the camera extrinsics of a pseudo view are computed as
@@
where $(\mathbf{R}_{1},\mathbf{R}_{2})$ represent the rotations of the sampled cameras, $(\mathbf{c}_{1},\mathbf{c}_{2})$ indicate the sampled camera centers, SLERP denotes the spherical linear interpolation [4], $\beta$ is a scalar that controls the interpolation, and $(\hat{\mathbf{R}},\hat{\mathbf{T}})$ stand for the camera parameters of the pseudo view. As illustrated in Fig. 3, for each pseudo view, we render an RGB image using the current $\mathbf{\Delta}$-model and store the renderings at different training steps in a buffer. We then compute a pixel-wise uncertainty map
$\mathbf{U}=\sqrt{\frac{1}{S}\sum_{i=1}^{S}(\mathbf{I}_{i}-\bar{\mathbf{I}})^{2}},$ (8)
of the same resolution as the rendered image, i.e., $\mathbf{U}\in\mathbb{R}^{H\times W}$, where $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size. This uncertainty estimation is carried out over all $M$ sampled pseudo views in parallel. In practice, to enhance robustness, we perform local smoothing over each uncertainty map, which yields
@@
where $k=5$ is the kernel size. Subsequently, we apply an uncertainty-aware perturbation to each of the Gaussians in the $\mathbf{\Delta}$-model as
@@
where $\hat{\mathcal{U}}^{t}$ is the set of the uncertainty maps computed from the buffers at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as
@@
with $\tau$ a predefined threshold, and $u_{ij}$ the uncertainty score of a pixel sampled from the set of pixels $\mathcal{P}$ overlapping with the 2D splats of $G_{\Delta}^{t}$. Please refer to the supplementary material for details on defining $\tau$ and identifying $\mathcal{P}$. In short, the reliability of each Gaussian is connected with the uncertainties of the renderings, and only the unreliable Gaussians, characterized by high uncertainty scores, are perturbed. In practice, we add the perturbation to the 3D positions, 3D rotations, scales, and opacities of the $\mathbf{\Delta}$-model. In particular, since 3D rotation is not continuous when expressed as a 3D matrix, we perturb the 6D continuous representation [51] of the rotation, which yields
$\hat{\mathbf{R}}^{t}_{\Delta}=f^{-1}(f(\mathbf{R}_{\Delta}^{t})+\delta^{R}_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})),\ \ \delta^{R}_{t}\in\mathbb{R}^{6},$ (12)
where $f(\cdot)$ denotes the mapping from rotation matrix to 6D representation. Fig. 4 illustrates the dynamic updates of the buffer throughout the training process. As the buffer is updated, the uncertainty map varies accordingly, resulting in diverse temporal samples from the $\mathbf{\Delta}$-model.
It is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to compute uncertainties directly in the Gaussian parameter space across different training steps. Our approach naturally handles this challenge as the statistics on 2D renderings provide a consistent ground to measure uncertainties, regardless of the varying number of Gaussians.
@@ 3.3 Self-Ensembling in 3DGS
Given the temporal samples around the $\mathbf{\Delta}$-model, the next step in our pipeline is to construct an ensemble from these samples. Concretely, we train a separate 3DGS model, named the $\mathbf{\Sigma}$-model, on the posed images without perturbations. To prevent the $\mathbf{\Sigma}$-model from being trapped at a suboptimal solution, we guide the training of the $\mathbf{\Sigma}$-model with a regularization [47] formulated as
@@
where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term [21], and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and $\mathbf{\Delta}$-model, respectively. We also utilize a co-pruning strategy [47] to enhance the regularization. Therefore, the $\mathbf{\Sigma}$-model aggregates information from different samples in the parameter space, serving as an ensemble of 3DGS models. The final loss function during training is defined as
@@
where $\gamma=1$ by default and $\mathcal{L}_{\text{RGB}}$ represents a photometric loss over the training views. During testing, we keep the $\mathbf{\Sigma}$-model for novel view synthesis. Notably, compared with the previous approach [47], our SE-GS inherently encodes a large number of 3DGS models without significant computational overhead, thereby enabling effective regularization efficiently. Furthermore, the regularization is carried out on pseudo views in a self-supervised manner, making it independent of additional information such as depth [24, 53].
@@ 4.1 Setup
Datasets. We conduct experiments on four datasets, i.e., LLFF [27], DTU [20], Mip-NeRF360 [3], and MVImgNet [45]. On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in [47], using the same training/testing splits and downsampling rates. As suggested in [47], we mask the background when assessing the quality of novel view synthesis on the DTU dataset. Since LLFF, DTU, and Mip-NeRF360 offer a limited number of scenarios, we extend our experiments to a large-scale dataset, i.e., MVImgNet. We randomly sample 50 scenes from this dataset and resize the longest side of each image to 512. The Gaussian parameters for all evaluated methods are initialized based on the point clouds obtained from COLMAP [36] to ensure a fair comparison. Note that COLMAP fails on certain DTU scenes in the sparse-view setting. In these cases, we instead randomly initialize the point cloud. Please refer to the supplementary material for more details on the setup.
Implementation details. We train our SE-GS for 10,000 iterations on the LLFF, DTU, and MVImgNet datasets, and for 30,000 iterations on Mip-NeRF360. In our uncertainty-aware perturbation mechanism, we employ $M=24$ image buffers with a buffer size of $S=3$ and perturb the $\mathbf{\Delta}$-model every 500 iterations. The noise $\delta_{t}$ in Eq. 10 is sampled from a normal distribution with a mean of 0, and the standard deviation is adaptively adjusted based on the magnitude of the Gaussian parameters. More details are provided in the supplementary material.
@@ 4.2 Quantitative Results
We provide the quantitative results of the evaluated approaches on the LLFF, DTU, Mip-NeRF360, and MVImgNet datasets in Table 1, Table 2, Table 3, and Table 4, respectively. The best, second-best, and third-best results in each column are highlighted in red, orange, and yellow, respectively. For MVSplat [7], we employ the pretrained model released by the authors and evaluate it on DTU. Notably, the per-scene optimized 3DGS methods significantly outperform MVSplat, highlighting the advantage of scene-level radiance field modeling for high-quality novel view synthesis. For these scene-level approaches, the performance of both NeRF and 3DGS methods consistently declines as the number of training views decreases, underscoring the challenge of few-shot novel view synthesis. In this context, our method surpasses all competitors across the four datasets. Note that in the sparse-view scenario, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
@@
Specifically, on the LLFF and DTU datasets, we train the methods using 3, 6, and 9 views. Our method achieves an improvement of $0.34$ in terms of PSNR with 3 views on LLFF. The improvement is larger on DTU, reaching a gain of $0.59$. Since the images in the Mip-NeRF360 dataset are captured from diverse viewpoints, we sample more training views, i.e., 12 and 24, following the setting used in [47]. On this dataset, our SE-GS achieves the highest PSNR and SSIM, as well as the lowest LPIPS and AVGE, demonstrating superior NVS quality under the challenges of sparse views with large-scale viewpoint variations. Additionally, to further assess the robustness across diverse scenarios, we conduct an experiment on 50 scenes sampled from the MVImgNet dataset. As reported in Table 4, our method yields consistently better results than previous state-of-the-art approaches when trained with 5, 7, and 10 views. Our SE-GS also demonstrates better compatibility with relatively dense views. For instance, the PSNR improvement increases from $0.26$ to $0.59$ as the number of training views varies from 5 to 10.
@@
The methods that leverage auxiliary data terms, such as DNGaussian [24] and FSGS [53], sometimes outperform 3DGS. However, in some cases, such as with 9 views on LLFF, the PSNR of FSGS becomes worse than 3DGS. Since the data acquired through off-the-shelf approaches might be unreliable, the quality of the NVS results is consequently affected. Compared with FSGS and DNGaussian, CoR-GS [47] exhibits better stability, outperforming 3DGS in more scenarios. This finding highlights the promising potential of regularization in the sparse-view setting. However, the improvement is still not consistent, as CoR-GS shows worse LPIPS than 3DGS when trained with 10 views on the MVImgNet dataset. By contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability. Our method is capable of avoiding specific local optima as the $\mathbf{\Sigma}$-model aggregates information from diverse temporal samples in the presented self-ensembling mechanism.
4.3 Qualitative Results
@@
Fig. 5 presents qualitative comparisons of 3DGS, CoR-GS, and our SE-GS, showing renderings from novel views. The models are trained on the DTU and MVImgNet datasets with 3 and 5 posed images, respectively. Our SE-GS produces fewer visual artifacts than the other methods, achieving better robustness against the challenge of few-shot training views. Moreover, our method captures finer details, particularly in areas with complex and repeated textures, such as the pineapple and watch strap. This demonstrates the effectiveness of our self-ensembling strategy in improving both the visual quality and stability of novel view neural rendering under sparse-view conditions.
@@
Figure 6: Comparison with CoR-GS. CoR-GS [47] is trained with different numbers of 3DGS models. The number varies from 2 to 8. The PSNR on LLFF with 3 training views is reported to evaluate the NVS quality. Iterations/s refers to the number of training iterations completed per second, indicating the training speed.
@@ 4.4 Analysis
To shed more light on the effectiveness and efficiency of our SE-GS, we perform a comprehensive analysis of the introduced self-ensembling mechanism. The analysis starts by comparing our approach with CoR-GS trained with varying numbers of Gaussian models. The detailed results on LLFF with 3 training views are shown in Fig. 6. As shown by Fig. 6(a), the PNSR of CoR-GS increases as more Gaussian models are trained. Since adding more Gaussian models enhances the regularization effect in CoR-GS, this observation highlights the potential of improving 3DGS in the sparse-view setting by strengthening the regularization process. The PSNR reaches a peak with 6 Gaussian models, indicating an upper bound of CoR-GS. Our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in more diverse samples of Gaussian models compared to CoR-GS. This leads to more effective regularization, as reflected by the better PSNR score.
Moreover, Fig. 6(b) shows the training speed measured as the number of training iterations completed per second. The speed of CoR-GS significantly drops as the number of Gaussian models increases. This limitation poses a challenge in scaling CoR-GS up to a large number of Gaussian models. By contrast, in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model. The temporal samples are generated through perturbations. We achieve a comparable training speed to CoR-GS with 2 models, showing that the perturbation process incurs a negligible additional cost. Note that only the $\mathbf{\Sigma}$-model is retained for NVS during inference, so the testing speed of SE-GS is the same as that of the original 3DGS. Consequently, our method is capable of efficiently enhancing the performance of 3DGS with sparse training views.
@@
Table 5: Effectiveness of the uncertainty-aware perturbations. Random indicates that we add random perturbations to all Gaussians in the $\mathbf{\Delta}$-model. Gradient-Aware denotes an alternative in which we perturb the Gaussians based on gradients. All methods are trained on LLFF with 3 views for 2000 iterations and the PSNR is utilized as the metric.
Finally, we assess the effectiveness of our uncertainty-aware perturbation strategy, comparing it with two alternatives. Specifically, as introduced in Sec. 3, a straightforward method for perturbing the $\mathbf{\Delta}$-model is to add random noise to the parameters of all Gaussians in the model. We denote this approach as Random. Building upon this baseline, an alternative approach, referred to as Gradient-Aware, identifies unreliable Gaussians by analyzing the gradients. Specifically, we replace the indicator function in Eq. 11 with one based on gradient magnitudes. Gaussians with gradient magnitudes exceeding a threshold are perturbed. In practice, we adopt the same threshold as used in density control. We train these two alternatives and our approaches on LLFF with 3 views for 2000 iterations, which challenges the methods in terms of convergence speed. As listed in Table 5, Random performs the worst since it leads to many invalid Gaussian samples, negatively impacting convergence speed. Note that our uncertainty-aware strategy aggregates information from multiple training steps, while the gradient-aware method can only utilize the gradients at the current training step. The improvement in PSNR shown in Table 5 demonstrates the superior effectiveness of our approach.